OpenStudy (elcna):

Help on this question? I'll fan and medal. :) What is the solution of the system? 2x-y=-7 4x-y=-4 *(-1.5,4) *(1.5,10) *(4,-1.5) *(-1.5,-2) I'm confused because I keep getting (3/2,10) and thats not an answer choice. Can someone explain what I'm doing wrong?

OpenStudy (blake57roger):

Rearrange the terms for either equation... Let's say the second equation. -y = -4x - 4 Substitute that for the first equation and solve for x... 2x - 4x - 4 = -7 -2x - 4 = -7 -2x = -3 x = 3/2 Finally... 2(3/2) - y = -7 3 - y = -7 -y = -10 y = 10 Hence, the solution is (3/2, 10) in terms of (x,y)

OpenStudy (blake57roger):

your right

OpenStudy (elcna):

Oh wow, it's just in decimal form on the answers! Haha I feel dumb, thankyou :)

OpenStudy (blake57roger):

welcame

OpenStudy (blake57roger):

welcome